OpenStudy (anonymous):

slope:1.6, y-intercept:7.5 choices: y=1.6x-7.5 y=-8/7x+7.5 y=1.6x+7.5 y=-1.6+7.5

OpenStudy (anonymous):

Again, this problem involves the use to the point slope equation

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y-7.5=1.6(x-0) then simplify the right side: y-7.5=1.6x-0, next as 7.5 to both sides of the equations, this results in : y=1.6x+7.5

OpenStudy (anonymous):

thank you